# Vojta’s abc conjecture for entire curves in toric varieties highly ramified over the boundary

### v1

@@ 4. The abc theorem for parabolic Riemann surfaces
Let $G$ be a non-constant polynomial in $\mathbb{C}[x_{1},\ldots,x_{n}]$ with no monomial factors and no repeated factors. Then, for any $\epsilon>0$, there exist a positive real number $c_{0}$ and a proper Zariski closed subset $Z$ of $\mathbb{A}^{n}(\mathbb{C})$ such that for all $n$-tuple $\mathbf{u}=(u_{1},\ldots,u_{n})$ of units on ${\mathcal{Y}}$ with $\mathbf{u}(\mathcal{Y})\not\subset Z$ and
(4.15) $\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}\geq c_{0}(\mathfrak{X}^{+}_{\sigma}(r)+\log r),$
we have
(4.16) $N_{G(\mathbf{u}),\mathcal{Y}}(0,r)-N^{(1)}_{G(\mathbf{u}),\mathcal{Y}}(0,r))\leq_{\rm exc}\epsilon\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}$
if $G(\mathbf{u})\neq 0$. Here, $c_{0}$ can be effectively bounded from above in terms of $\epsilon$, $n$, and the degree of $G$. Moreover, the exceptional set $Z$ can be expressed as the zero locus of a finite set $\Sigma\subset\mathbb{C}[x_{1},\ldots,x_{n}]$ with the following properties: (Z1) $\Sigma$ depends on $\epsilon$ and $G$, but not $\mathcal{Y}$ and can be determined explicitly, (Z2) $|\Sigma|$ and the degree of each polynomial in $\Sigma$ can be effectively bounded from above in terms of $\epsilon$, $n$, and the degree of $G$.
@@
We follow the proof of [7, Theorem 4]. Let $u_{i},1\leq i\leq n$, be units on ${\mathcal{Y}}$. We apply Theorem 19 through several reduction steps to derive a contradiction. Assume that (4.16) does not hold and let $c_{0}\geq\tilde{c}_{1}$, where $\tilde{c}_{1}$ is the constant from (4.5) in Theorem 19. Then, there exists a positive integer $M_{1}$, depending only on $\epsilon$, $n$, and the degree of $G$, but independent of $\mathbf{u}$, such that
(4.17) $T_{u_{1}^{m_{1}}\cdots u_{n}^{m_{n}},\mathcal{Y}}(r)\leq_{\rm exc}\tilde{c}_{2}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O\left(\max_{1\leq j\leq n}\{\log T_{u_{j},{\mathcal{Y}}}(r)\}\right)$
for a non-trivial $n$-tuple $(m_{1},\ldots,m_{n})$ of integers with $\sum|m_{i}|\leq M_{1}$. Thus, we have
(4.18) $\lambda_{1}:={u_{1}}^{m_{1}}\cdots{u_{n}}^{m_{n}}\in\mathbb{K}_{\mathbf{u}}.$
@@
For the particular change of variables in (4.22), (4.23), and (4.24) (that depends on the matrix $A_{1}$), we apply the Lemma 16 in [7] to find a nonconstant polynomial $H_{1}^{\prime}\in\mathbb{C}[x_{1},\ldots,x_{n}]$ such that $G_{1}(\lambda_{1},X_{1,2},\ldots,X_{1,n})$ has neither monomial nor repeated irreducible factors if $H_{1}^{\prime}(u_{1},\ldots,u_{n})\not\equiv 0$. It is important to note that the polynomial $H_{1}^{\prime}$ is not only independent of $\mathbf{u}$, but also independent of $\mathcal{Y}$. We now take $H_{1}$ to be the product of all such $H_{1}^{\prime}$ where $A_{1}$ ranges over the finitely many elements of $\operatorname{GL}_{n}(\mathbb{Z})$ with $\|A_{1}\|_{\infty}\leq M_{1}+n$. From Lemma 16 in [7]  we know that $\deg H_{1}$ depends only on $\epsilon$, $n$ and $\deg G$.
@@
(4.25) $\begin{split}N_{G(\mathbf{u}),\mathcal{Y}}(0,r)-N_{G(\mathbf{u}),\mathcal{Y}}^{(1)}(0,r)=N_{G_{1}(\lambda_{1},\beta_{1,2},\ldots,\beta_{1,n}),\mathcal{Y}}(0,r)-N_{G_{1}(\lambda_{1},\beta_{1,2},\ldots,\beta_{1,n}),\mathcal{Y}}^{(1)}(0,r).\end{split}$
On the other hand, by taking $\tilde{c}_{1}\geq 2\tilde{c}_{2}\epsilon^{-1}$, we deduce from (4.15) and (4.17) that
(4.26) $T_{{u_{1}}^{m_{1}}\cdots{u_{n}}^{m_{n}},\mathcal{Y}}(r)\ \leq_{\rm exc}\epsilon\max_{1\leq j\leq n}\{T_{u_{j},{\mathcal{Y}}}(r)\}.$
Then, by (4.22), (4.23) and (4.26), we have:
$\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}={\rm O}\left(\max\{T_{\lambda_{1},\mathcal{Y}}(r),T_{\beta_{1,2},\mathcal{Y}}(r),\ldots,T_{\beta_{1,n},\mathcal{Y}}(r)\}\right)={\rm O}\left(\max_{2\leq i\leq n}\{T_{\beta_{1,i},\mathcal{Y}}(r)\}\right).$
More precisely, we find positive real numbers $b_{1,1}$ and $b_{1,2}$ such that
(4.27) $b_{1,1}\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}\leq\max_{2\leq i\leq n}\{T_{\beta_{1,i},\mathcal{Y}}(r)\},\quad\text{and }\quad\max_{2\leq i\leq n}\{T_{\beta_{1,i},\mathcal{Y}}(r)\}\leq b_{1,2}\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}.$
In conclusion, at the end of this step we have that
$\max_{2\leq i\leq n}\{T_{\beta_{1,i},\mathcal{Y}}(r)\}\geq b_{1,1}c_{0}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)$
holds, and the inequality
(4.28) $N_{G_{1}(\lambda_{1},\beta_{1,2},\ldots,\beta_{1,n})}(0,r)-N_{G_{1}(\lambda_{1},\beta_{1,2},\ldots,\beta_{1,n})}^{(1)}(0,r)$ $\leq_{\rm exc}\frac{\epsilon}{b_{1,2}}\max_{2\leq i\leq n}\{T_{\beta_{1,i},\mathcal{Y}}(r)\}$
@@
The above is the first step. There are $(n-1)$ many steps in total as in the proof of [7, Theorem 4]. Let $12\leq s\leq n-$ 1 and suppose that we have completed the step $s-1$. This includes the construction of $H_{s-1}\in{\mathbb{C}}[x_{1},\ldots,x_{n}]$ with degree depends on $\epsilon$, $n$ and $\deg G$ only. The last one is Step $(n-1)$ resulting in $H_{n-1}\in{\mathbb{C}}[x_{1},\ldots,x_{n}]$. We now define $H=H_{1}\cdots H_{n-1}$. Then $\deg H$ depends only on $\epsilon$, $n$ and $\deg G$, since each $H_{i}$ does so. Moreover, the polynomial $H$ is not only independent of $\mathbf{u}$, but also independent of $\mathcal{Y}$.
Suppose $H(u_{1},\ldots,u_{n})\not\equiv 0$. Assume we go through all the $(n-1)$ steps to get the polynomial
@@
such that its degree can be bounded explicitly in terms of $M_{n-1},\ldots,M_{1}$, $n$, and $\deg(G)$. At the end of the step $(n-1)$, we have that $\beta:=\beta_{n-1,n}$, which is a (non-constant) unit on ${\mathcal{Y}}$, such that
(4.29) $T_{\beta,\mathcal{Y}}(r)\geq b_{n-1,1}c_{0}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)$
holds and the inequality
(4.30) $N_{P(\beta)}(0,r)-N^{(1)}_{P(\beta)}(0,r)\leq_{\rm exc}\frac{\epsilon}{b_{n-1,2}}T_{\beta,\mathcal{Y}}(r)$
@@
(4.31) $N_{\gcd,{\mathcal{Y}}}(P(\beta),D_{\beta}(P)(\beta),r)\geq N_{P(\beta)}(0,r)-N^{(1)}_{P(\beta)}(0,r).$
@@
(4.32) $N_{\gcd,{\mathcal{Y}}}(P(\beta),D_{\beta}(P)(\beta),r)\leq C_{1}(T_{P,{\mathcal{Y}}}(r)+T_{D_{\beta}(P),{\mathcal{Y}}}(r)).$
Similar to the arguments for (4.9), we find a positive constant $C_{2}$ independent of $\beta$ such that
(4.33) $T_{P,{\mathcal{Y}}}(r)+T_{D_{\beta}(P),{\mathcal{Y}}}(r)\leq_{\rm exc}C_{2}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(\log T_{\beta,\mathcal{Y}}(r)).$
Then by (4.29), (4.31), (4.32), and (4.33), we have
(4.34) $N_{P(\beta)}(0,r)-N^{(1)}_{P(\beta)}(0,r)\leq_{\rm exc}\frac{C_{1}C_{2}}{b_{n-1,1}c_{0}}T_{\beta,\mathcal{Y}}(r)+O(\log T_{\beta,\mathcal{Y}}(r)),$
which contradicts (4.30) by taking $c_{0}\geq 2C_{1}C_{2}b_{n-1,2}b_{n-1,1}^{-1}\epsilon^{-1}$. ∎
@@
Let $G=\sum_{\mathbf{i}\in I_{G}}\alpha_{\mathbf{i}}{\mathbf{x}}^{\mathbf{i}}\in$ be a non-constant polynomial in $\mathbb{C}[x_{1},\cdots,x_{n}]$. Assume that $G(0,\ldots,0)\neq 0$ and $\deg_{x_{i}}G=\deg G=d$ for $1\leq i\leq n$. Let $Z_{1}$ be the Zariski closed subset that is the union of hypersurfaces of $\mathbb{A}^{n}$ of the form $\sum_{\mathbf{i}\in J}\alpha_{\mathbf{i}}{\mathbf{x}}^{\mathbf{i}}=0$, where $J$ is a non-empty subset of $I_{G}$. Let $u_{1},\ldots,u_{n}$ be the units on $\mathcal{Y}$ such that the image of ${\bf u}:=(u_{1},\ldots,u_{n}):\mathcal{Y}\to(\mathbb{C}^{*})^{n}$ is not contained in $Z_{1}$. Then
(4.35) $N_{G({\bf u}),\mathcal{Y}}(0,r)\geq_{\operatorname{exc}}(\deg G-\epsilon)\cdot\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}-c(\mathfrak{X}^{+}_{\sigma}(r)+\log r),$
@@
where $\alpha_{\mathbf{i}_{i}}\neq 0$ for $0\leq i\leq n$ and $I=\{\mathbf{i}_{0}:=(0,0,\ldots,0),\mathbf{i}_{1}:=(d,0,\ldots,0),\ldots,\mathbf{i}_{n}:=(0,\ldots,0,d)\}$. If ${\mathbf{u}}({\mathbb{C}})$ is not contained in $Z_{1}$, we may use Theorem 12 to show that
$(d-\epsilon)\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}\leq N_{G(\mathbf{u}),\mathcal{Y}}(0,r)+c(\mathfrak{X}^{+}_{\sigma}(r)+\log r),$
where $c=\frac{1}{2}\binom{n+d}{n}(\binom{n+d}{n}+1)$. ∎
@@
Let $G$ be a non-constant polynomial in $\mathbb{C}[x_{1},\ldots,x_{n}]$ with no monomial factors and no repeated factors. Then, for any $\epsilon>0$, there exist a positive real number $c_{0}$ and a proper Zariski closed subset $W$ of $\mathbb{A}^{n}(\mathbb{C})$ such that for all holomorphic maps $\mathbf{u}=(u_{1},\ldots,u_{n}):\mathcal{Y}\to(\mathbb{C}^{*})^{n}$ with $\mathbf{u}(\mathcal{Y})\not\subset W$ and satisfying
(4.37) $\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}\geq c_{0}(\mathfrak{X}^{+}_{\sigma}(r)+\log r),$
we have the following two statements hold
@@
    $N_{G(\mathbf{u}),\mathcal{Y}}(0,r)-N^{(1)}_{G(\mathbf{u}),\mathcal{Y}}(0,r))\leq_{\rm exc}\epsilon\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}$.
@@
    $N^{(1)}_{G(\mathbf{u}),\mathcal{Y}}(0,r))\geq_{\rm exc}(\deg G-\epsilon)\cdot\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}.$
Here, $c_{0}$ can be effectively bounded from above in terms of $\epsilon$, $n$, and the degree of $G$. Moreover, the exceptional set $W$ can be expressed as the zero locus of a finite set $\Sigma\subset K[x_{1},\ldots,x_{n}]$ with the following properties: (Z1) $\Sigma$ depends on $\epsilon$ and $G$ (not on $\mathcal{Y}$) and can be determined explicitly, (Z2) $|\Sigma|$ and the degree of each polynomial in $\Sigma$ can be effectively bounded from above in terms of $\epsilon$, $n$, and the degree of $G$.
@@
The proof follows from Theorem 20 and Theorem 6.4. Let $W=Z\cup Z_{1}\cup[G=0]$, where $Z$ and $Z_{1}$ are the exceptional set from Theorem 20 and Theorem 6.4 for $\frac{\epsilon}{2}$ respectively. Let $\mathbf{u}=(u_{1},\ldots,u_{n}):\mathcal{Y}\to(\mathbb{C}^{*})^{n}$ with $\mathbf{u}(\mathcal{Y})\not\subset W$. Theorem 20 implies that
$N_{G(\mathbf{u}),\mathcal{Y}}(0,r)-N^{(1)}_{G(\mathbf{u}),\mathcal{Y}}(0,r))\leq\frac{\epsilon}{2}\max_{1\leq i\leq n}\{T_{u_{i}}(r)\}.$
Together with (4.35) in Theorem 6.4, we derive
(4.38) $N^{(1)}_{G(\mathbf{u}),\mathcal{Y}}(0,r))\geq_{\operatorname{exc}}\left(\deg G-\frac{\epsilon}{2}\right)\max_{1\leq i\leq n}\{T_{u_{i}}(r)\}-c(\mathfrak{X}^{+}_{\sigma}(r)+\log r).$
By replacing the constant $c_{0}$ in (4.37) by $c_{1}=\max\{c_{0},2c\epsilon^{-1}\}$, we may derive from (4.38) that
$N^{(1)}_{G(\mathbf{u}),\mathcal{Y}}(0,r))\geq_{\operatorname{exc}}(\deg G-\epsilon)\cdot\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}.$

### v2

@@ 4. The abc theorem for parabolic Riemann surfaces
Let $G$ be a non-constant polynomial in $\mathbb{C}[x_{1},\ldots,x_{n}]$ with no monomial factors and no repeated factors. Then, for any $\epsilon>0$, there exist a positive real number $c_{0}$, a proper Zariski closed subset $Z$ of $\mathbb{A}^{n}(\mathbb{C})$ expressible as the zero locus of a finite set $\Sigma\subset\mathbb{C}[x_{1},\ldots,x_{n}]$ with the following property: Let ${\bf u}=(u_{1},\dots,u_{n})$, where each $u_{i}$, $1\leq i\leq n$, is meromorphic function on $\mathbb{C}$}, with $H({\mathbf{u}})\neq 0$ for all $H\in\Sigma$ and $G(\mathbf{u})\neq 0$. Let $\mathcal{Y}$$\subset{\mathbb{C}}$ be given as above with $\sigma$ being associated to the discs $\{{\mathbb{D}}(a_{j},2r_{j})\}$. Assume that $u_{1},\dots,u_{n}$ are units on $\mathcal{Y}$}. Then, by shrinking $r_{j}$ small enough (depending on $\bf u$} and $G$), we have
(4.12) $N_{G(\mathbf{u}),\sigma}(0,r)-N^{(1)}_{G(\mathbf{u}),\sigma}(0,r))\leq_{\rm exc}\epsilon\max_{1\leq i\leq n}\{T_{u_{i}}(r)\}+c_{0}(\mathfrak{X}^{+}_{\sigma}(r)+\log r).$
Here, $c_{0}$ can be effectively bounded from above in terms of $\epsilon$, $n$, and the degree of $G$. Moreover, the finite set $\Sigma\subset\mathbb{C}[x_{1},\ldots,x_{n}]$ satisfies: (Z1) $\Sigma$ depends on $\epsilon$ and $G$, but not on $\bf u$} or $\mathcal{Y}$, and can be determined explicitly, (Z2) $|\Sigma|$ and the degree of each polynomial in $\Sigma$ can be effectively bounded from above in terms of $\epsilon$, $n$, and the degree of $G$.
@@
We follow the proof of [7, Theorem 4]. We apply Theorem 23 through several reduction steps to derive a contradiction. The constant $c_{0}$ will be determined in the course of the proof. We also remark that, since the coefficients of $G$ lie in $\mathbb{C}$, the assumption that at least one coefficient in the expansion of $G$ is equal to $1$ imposes no restriction. By Theorem 23, after shrinking $r_{j}$, we have either
(4.13) $T^{M}_{u_{1}^{m_{1}}\cdots u_{n}^{m_{n}},\sigma}(r)\leq_{\rm exc}\tilde{c}_{1}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(\max_{1\leq j\leq n}\{\log T_{u_{j}}(r)\})$
or
(4.14) $N_{G(\mathbf{u}),\sigma}(0,r)-N^{(1)}_{G(\mathbf{u}),\sigma}(0,r)\leq_{\rm exc}\epsilon\max_{1\leq j\leq n}\{T_{u_{j}}(r)\}+\tilde{c}_{2}(\mathfrak{X}^{+}_{\sigma}(r)+\log r).$
Let $c_{0}\geq\tilde{c}_{2}$. If (4.14) holds, then we are done. Assume it is not, then (4.13) holds for a non-trivial $n$-tuple $(m_{1},\ldots,m_{n})$ of integers with $\sum|m_{i}|\leq M_{1}$, where $M_{1}$, depending only on $\epsilon$, $n$, and the degree of $G$, but independent of $\mathbf{u}$. Thus, we have
(4.15) $\lambda_{1}:={u_{1}}^{m_{1}}\cdots{u_{n}}^{m_{n}}\in\mathbb{K}_{\mathbf{u}}.\$
@@
For the particular change of variables in (4.19), (4.20), and (4.21) (that depends on the matrix $A_{1}$), we apply the Lemma 16 in [7] to find a nonconstant polynomial $\widetilde{H}_{1}\in\mathbb{C}[x_{1},\ldots,x_{n}]$ such that $G_{1}(\lambda_{1},X_{1,2},\ldots,X_{1,n})$ has neither monomial nor repeated irreducible factors if $\widetilde{H}_{1}(u_{1},\ldots,u_{n})\not\equiv 0$. It is important to note that the polynomial $\widetilde{H}_{1}$ is not only independent of $\mathbf{u}$, but also independent of $\mathcal{Y}$. We now take $H_{1}$ to be the product of all such $\widetilde{H}_{1}$ where $A_{1}$ ranges over the finitely many elements of $\operatorname{GL}_{n}(\mathbb{Z})$ with $\|A_{1}\|_{\infty}\leq M_{1}+n$. From Lemma 16 in [7]  we know that $\deg H_{1}$ depends only on $\epsilon$, $n$ and $\deg G$.
@@
(4.22) $\begin{split}N_{G(\mathbf{u}),\sigma}(0,r)-N_{G(\mathbf{u}),\sigma}^{(1)}(0,r)=N_{G_{1}(\lambda_{1},\beta_{1,2},\ldots,\beta_{1,n}),\sigma}(0,r)-N_{G_{1}(\lambda_{1},\beta_{1,2},\ldots,\beta_{1,n}),\sigma}^{(1)}(0,r).\end{split}$
By (4.19) and (4.20), we have:
(4.23) $\max_{1\leq i\leq n}\{T_{u_{i}}(r)\}={\rm O}\left(\max\{T_{\lambda_{1}}(r),T_{\beta_{1,2}}(r),\ldots,T_{\beta_{1,n}}(r)\}\right).$
By (4.16), there are only a finitely many choice of $\beta_{1,t}$, $2\leq t\leq n$. Hence, by Proposition 12, after possibly shrinking $r_{j}$ (depending on $u_{i}$ ), we may assume that
$\bigl|T_{u_{i}}(r)-T^{M}_{u_{i},\sigma}(r)\bigr|,\quad\bigl|T_{\beta_{1,t}}(r)-T^{M}_{\beta_{1,t},\sigma}(r)\bigr|,\quad\text{and}\quad\bigl|T_{\lambda_{1}}(r)-T^{M}_{\lambda_{1},\sigma}(r)\bigr|$
are all bounded by $\log r+O(1)$ for $1\leq i\leq n$ and $2\leq t\leq n$. Then by (4.23) and (4.13), we find positive real numbers $b_{1,1}$ and $b_{1,2}$ such that
$b_{1,1}\max_{1\leq i\leq n}\{T_{u_{i}}(r)\}\leq_{\rm exc}\max_{2\leq t\leq n}\{T_{\beta_{1,t}}(r)\}+\tilde{c}_{1}(\mathfrak{X}^{+}_{\sigma}(r)+\log r),\quad\text{and}$
$\max_{2\leq t\leq n}\{T_{\beta_{1,t}}(r)\}\leq b_{1,2}\left(\max_{1\leq i\leq n}\{T_{u_{i}}(r)\}+\log r\right).$
In conclusion, at the end of this step we have that, after possibly shrinking $r_{j}$, the inequality
(4.24) $N_{G_{1}(\lambda_{1},\beta_{1,2},\ldots,\beta_{1,n}),\sigma}(0,r)-N_{G_{1}(\lambda_{1},\beta_{1,2},\ldots,\beta_{1,n}),\sigma}^{(1)}(0,r)$ $\leq_{\rm exc}\frac{\epsilon}{b_{1,2}}\max_{2\leq i\leq n}\{T_{\beta_{1,i}}(r)\}+c_{0}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)$
@@
The above is the first step. The next step is to apply Theorem 23 to $G_{1}$ and $\bf u$}. Note that, to do so, we need to further shrink $r_{j}$ according to $G_{1}$. There are (n1)(n-1) many steps in total as in the proof of [7, Theorem 4]. Let $2\leq s\leq n-1$ and suppose that we have completed the step s1s-1. This includes the construction of $H_{s-1}\in{\mathbb{C}}[x_{1},\ldots,x_{n}]$ with degree depends on $\epsilon$, $n$ and $\deg G$ only. The last one is Step (n1)(n-1) resulting in $H_{n-1}\in{\mathbb{C}}[x_{1},\ldots,x_{n}]$. We now define $H=H_{1}\cdots H_{n-1}$. Then $\deg H$ depends only on $\epsilon$, $n$ and $\deg G$, since each $H_{i}$ does so. Moreover, the polynomial $H$ is not only independent of $\mathbf{u}$, but also independent of $\mathcal{Y}$.
Suppose $H(u_{1},\ldots,u_{n})\not\equiv 0$. Assume we go through all the (n1)(n-1) steps to get the polynomial
@@
such that its degree can be bounded explicitly in terms of $M_{n-1},\ldots,M_{1}$, $n$, and $\deg(G)$. At the end of the step (n1)(n-1), we have that $\beta:=\beta_{n-1,n}$, which is a (non-constant) unit on $\mathcal{Y}$}, such that the inequality, after further shrinking $r_{j}$ according to $G,G_{1},\dots,G_{n-1}$,
(4.25) $N_{P(\beta),\sigma}(0,r)-N^{(1)}_{P(\beta),\sigma}(0,r)\leq_{\rm exc}\frac{\epsilon}{b_{n-1,2}}T_{\beta}(r)+c_{0}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)$
@@
(4.26) $N_{\gcd,\sigma}(P(\beta),D_{\beta}(P)(\beta),r)\geq N_{P(\beta),\sigma}(0,r)-N^{(1)}_{P(\beta),\sigma}(0,r).$
@@
(4.27) $N_{\gcd,\sigma}(P(\beta),D_{\beta}(P)(\beta),r)\leq C_{1}(T_{P,\sigma}(r)+T_{D_{\beta}(P),\sigma}(r)).$
Similar to the arguments for (4), after further shrinking $r_{j}$ according to $P$, we find a positive constant $C_{2}$ independent of $\beta$ such that
(4.28) $T_{P,\sigma}(r)+T_{D_{\beta}(P),\sigma}(r)\leq_{\rm exc}C_{2}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(\log T_{\beta}(r)).$
Hence,
(4.29) $N_{P(\beta),\sigma}(0,r)-N^{(1)}_{P(\beta),\sigma}(0,r)\leq_{\rm exc}C_{1}C_{2}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(\log T_{\beta}(r)),$
which contradicts that fact that (4.25) fails to hold by taking $c_{0}\geq C_{1}C_{2}$. ∎
@@
Let $\mathcal{Y}$$\subset{\mathbb{C}}$ be given as above with $\sigma$ being associated to the discs $\{{\mathbb{D}}(a_{j},2r_{j})\}$. Let $G=\sum_{\mathbf{i}\in I_{G}}\alpha_{\mathbf{i}}{\mathbf{x}}^{\mathbf{i}}$ be a non-constant polynomial in $\mathbb{C}[x_{1},\cdots,x_{n}]$. Assume that $G(0,\ldots,0)\neq 0$ and $\deg_{x_{i}}G=\deg G=d$ for $1\leq i\leq n$. Let $f_{0},\dots,f_{n}$ be entire functions without common zero which are units on $\mathcal{Y}$. Let $u_{i}:=f_{i}/f_{0}$, $1\leq i\leq n$, and ${\bf u}=(u_{1},\dots,u_{n})$. Assume that $\sum_{\mathbf{i}\in J}\alpha_{\mathbf{i}}{\mathbf{u}}^{\mathbf{i}}\neq 0$ for any non-empty subset $J$ of $I_{G}$. Then, for every $\epsilon>0$, by shrinking $r_{j}$ small enough (depending on $\epsilon$, $\bf u$} and $G$), we have
(4.30) $N_{G({\bf u}),\sigma}(0,r)\geq_{\operatorname{exc}}(\deg G-\epsilon)\cdot\max_{1\leq i\leq n}\{T_{u_{i}}(r)\}-c(\mathfrak{X}^{+}_{\sigma}(r)+\log r),$
@@
where $\alpha_{\mathbf{i}_{i}}\neq 0$ for $0\leq i\leq n$ and $I=\{\mathbf{i}_{0}:=(0,0,\ldots,0),\mathbf{i}_{1}:=(d,0,\ldots,0),\ldots,\mathbf{i}_{n}:=(0,\ldots,0,d)\}$. Since $\sum_{\mathbf{i}\in J}\alpha_{\mathbf{i}}{\mathbf{u}}^{\mathbf{i}}\neq 0$, for any non-empty subset $J$ of $I_{G}$, we may use Theorem 16 to show that
$d\max_{1\leq i\leq n}\{T^{M}_{u_{i},\sigma}(r)\}\leq N_{G(\mathbf{u}),\sigma}(0,r)+c(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+c^{+}_{h_{\mathbf{u}},\sigma}(r)+O(\max_{1\leq i\leq n}\log T_{u_{i}}(r)),$
where $c=\frac{1}{2}\binom{n+d}{n}(\binom{n+d}{n}+1)$, and $h_{\mathbf{u}}$ is a meromorphic function depending on $\bf u$}. By Lemma 10, by shrinking $r_{j}$ small enough if necessary, we can make $c^{+}_{h_{\mathbf{u}},\sigma}(r)\leq O(1)$. In addition, we can shrink $r_{j}$ small enough such that (3.15) and the Proposition 12 for $u_{i},1\leq i\leq n$, hold, i.e., $|T^{M}_{u_{i},\sigma}(r)-T_{u_{i}}(r)|\leq\log r+O(1)$. This proves the theorem. ∎
@@
Let $\mathcal{Y}$$\subset{\mathbb{C}}$ be given with $\sigma$ being associated to the discs $\{{\mathbb{D}}(a_{j},2r_{j})\}$. Let $G$ be a non-constant polynomial in $\mathbb{C}[x_{1},\ldots,x_{n}]$ with no monomial factors and no repeated factors. Then, for any $\epsilon>0$, there exist a positive real number $c_{0}$, a proper Zariski closed subset $W$ of $\mathbb{A}^{n}(\mathbb{C})$ expressible as the zero locus of a finite set $\Sigma\subset\mathbb{C}[x_{1},\ldots,x_{n}]$ with the following property: Let $f_{0},\dots,f_{n}$ be entire functions without common zeros, which are units on $\mathcal{Y}$. For $1\leq i\leq n$, set $u_{i}:=f_{i}/f_{0}$ and write ${\bf u}=(u_{1},\dots,u_{n})$. Assume that $H({\bf u})\neq 0$ for all $H\in\Sigma$. Then, by shrinking $r_{j}$ small enough (depending on $\epsilon$, $\bf u$} and $G$), we have
@@
    $N_{G(\mathbf{u}),\sigma}(0,r)-N^{(1)}_{G(\mathbf{u}),\sigma}(0,r))\leq_{\rm exc}\epsilon\max_{1\leq i\leq n}\{T_{u_{i}}(r)\}+c_{0}(\mathfrak{X}^{+}_{\sigma}(r)+\log r).$
@@
    $(\deg G-\epsilon)\cdot\max_{1\leq i\leq n}\{T_{u_{i}}(r)\}\leq_{\rm exc}N^{(1)}_{G(\mathbf{u}),\sigma}(0,r))+c_{1}(\mathfrak{X}^{+}_{\sigma}(r)+\log r).$
Here, $c_{0}$ and $c_{1}$ can be effectively bounded from above in terms of $\epsilon$, $n$, and the degree of $G$. Moreover, the finite set $\Sigma\subset\mathbb{C}[x_{1},\ldots,x_{n}]$ satisfies: (Z1) $\Sigma$ depends on $\epsilon$ and $G$, but not on $\bf u$} or $\mathcal{Y}$, and can be determined explicitly, (Z2) $|\Sigma|$ and the degree of each polynomial in $\Sigma$ can be effectively bounded from above in terms of $\epsilon$, $n$, and the degree of $G$.
@@
The proof follows from Theorem 24 and Theorem 25. Let $W=Z\cup Z_{1}\cup[G=0]$, where $Z$ and $Z_{1}$ are the exceptional sets arising from Theorem 24 and Theorem 25 for $\frac{\epsilon}{2}$, respectively. We express the zero locus of $W$ as the vanishing set of a finite collection $\Sigma\subset\mathbb{C}[x_{1},\ldots,x_{n}]$, and assume that $H({\bf u})\neq 0$ for all $H\in\Sigma$. By Theorem 24, we obtain
$N_{G(\mathbf{u}),\sigma}(0,r)-N^{(1)}_{G(\mathbf{u}),\sigma}(0,r))\leq\frac{\epsilon}{2}\max_{1\leq i\leq n}\{T_{u_{i}}(r)\}+c_{0}(\mathfrak{X}^{+}_{\sigma}(r)+\log r).$
On the other hand, from (4.30) in Theorem 25,
$N_{G({\bf u}),\sigma}(0,r)\geq_{\operatorname{exc}}(\deg G-\epsilon)\cdot\max_{1\leq i\leq n}\{T_{u_{i}}(r)\}-c(\mathfrak{X}^{+}_{\sigma}(r)+\log r).$
Thus we derive
(4.32) $N^{(1)}_{G(\mathbf{u}),\sigma}(0,r))\geq_{\operatorname{exc}}\left(\deg G-\frac{\epsilon}{2}\right)\max_{1\leq i\leq n}\{T_{u_{i}}(r)\}-c(\mathfrak{X}^{+}_{\sigma}(r)+\log r).$
Let $c_{1}=c_{0}+c$, then
$N^{(1)}_{G(\mathbf{u}),\sigma}(0,r))\geq_{\operatorname{exc}}(\deg G-\epsilon)\cdot\max_{1\leq i\leq n}\{T_{u_{i}}(r)\}+c_{1}(\mathfrak{X}^{+}_{\sigma}(r)+\log r).$